OpenStudy (gabylovesyou):

@misty1212 What is the discontinuity and zero of the function f(x) = the quantity of 2x squared plus 5x minus 12, all over x plus 4?

OpenStudy (anonymous):

f(x) = (2x^2 + 5x - 12)/(x + 4) as given, is not defined when the denominator is 0 ie when x + 4 = 0 or x = -4. However, 2x^2 + 5x - 12 = (x + 4)*(2x - 3) so that the function can be simplified to (x + 4)*(2x - 3)/(x + 4) = 2x - 3 This has a zero when 2x - 3 = 0 ie 2x = 3 or x = 1.5

OpenStudy (misty1212):

\[f(x)=\frac{2x^2+5x-12}{x+4}\]right?

OpenStudy (gabylovesyou):

\[f(x) = \frac{ 2x^2 + 5x - 12 }{ x + 4 }\]

OpenStudy (gabylovesyou):

yes :)

OpenStudy (misty1212):

yeah that is what i thought

OpenStudy (misty1212):

set the deniminator equal to zero \[x+4=0\] and solve for \(x\) \[x=-4\]

OpenStudy (misty1212):

it is not continuous at \(x=-4\)

OpenStudy (misty1212):

same as what @muneeb180 said above

OpenStudy (gabylovesyou):

Discontinuity at (−4, −11), zero at (three halves , 0) Discontinuity at (−4, −11), zero at (negative three halves , 0) Discontinuity at (4, 5), zero at ( three halves , 0) Discontinuity at (4, 5), zero at (negative three halves , 0)

OpenStudy (gabylovesyou):

these are my choices..

OpenStudy (misty1212):

as for the zero, we have \[\frac{2x^2+5x-12}{x+4}=\frac{(x+4)(2x-3)}{x+4}=2x-3\]

OpenStudy (misty1212):

discontinuity find by putting \(-4\) in to \(2x-3\) and get \[2(-4)-3=-8-3=-11\] so it is at \((-4,-11)\)

OpenStudy (shadowlegendx):

Ignore him

OpenStudy (misty1212):

lol that child has never seen any in the flesh i am sure, only on porn sites

OpenStudy (gabylovesyou):

so we know its or A or B.

OpenStudy (anonymous):

yeah and zeros are at \[2x-3=0\\ 2x=3\\ x=\frac{3}{2}\]

OpenStudy (gabylovesyou):

lmfao so A,.. ok thanks ^-^

OpenStudy (misty1212):

what @satellite73 said

OpenStudy (misty1212):

go with A

OpenStudy (gabylovesyou):

:) thanks.

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]